If the vertices of square A are the midpoints of the edges of square B, then what is the ratio of the area of square A to square B?
Answer by edjones(8007)

Let s=side of the larger triangle.
s^2 = area of larger triangle
A hypotenuse of a triangle formed by .5s and .5s = a side of the smaller triangle.
2a^2=c^2
2(.5s)^2=.5s^2 Area of the smaller square and also square of the hypotenuse.
.5s^2/s^2=.5 ratio of the area of square A to square B.
.
Ed